30. A researcher reports an F-ratio with df 2, 18 from an independent-measures research study. Based on the df values, how mnay treatments were compared in the study and what was the total number of subjects participating in the study? a. 2 treatments and 19 subjects b. 2 treatments and 20 subjects c. 3 treatments and 21 subjects d. 3 treatments and 22 subjects 31. Two variables are said to int
Answer #1

Answer:

c. 3 treatments and 21 subjects

Solution:

We are given,

df1 = 2

that is,

df1 = Between degrees of freedom = number of treatments – 1

Number of treatments = df1 + 1 = 2 + 1 = 3

Now, we have

df2 = within df = 18

Total number of subjects = df1 + df2 + 1 = 2 + 18 + 1 = 21